# Representing in Low Rank I: conjugacy, topological and homological aspects

### v1

@@ 1. Introduction
This problem is highly dependent on the choice of the ground field $F$. For example if $F={\mathbb{C}}$, then $FG\cong FH$ for $G$ and $H$ abelian exactly when $|G|=|H|$. On the other hand, if $F={\mathbb{Q}}$ then the group algebra determines all the structure constants of an abelian group by a theorem of Perlis–Walker [PW]. One reason that $\mathbb{Q}$} detects more is the theorem of Wedderburn–Artin, which says that there is an isomorphism of $F$-algebras
@@
where $D_{i}$ are finite-dimensional division $F$-algebras. In particular, the Brauer group ${\rm Br}(F)$, i.e. the division algebras over $F$, has an influence. One way to view this is, the ‘smaller’ $F$, the ‘richer’ its Brauer group. Nevertheless, Dade [Dade] constructed non-isomorphic groups $G$ and $H$ with isomorphic group algebras over any field of characteristic 0, or in other words such that ${\mathbb{Q}}G\cong{\mathbb{Q}}H$. For a long time, it was believed that over the integers $\mathbb{Z}$} (equivalently over all group rings $RG$), a group $G$ is uniquely determined by its integral group ring ${\mathbb{Z}}G$. However, Hertweck constructed a counterexample to this [Hertweck]. Nevertheless, if ${\mathbb{Z}}G\cong{\mathbb{Z}}H$, then $G$ and $H$ are cocycle twists of each under a Cech-type non-abelian cohomology theory constructed by Roggenkamp–Kimmerle [KR]. Thus, although Hertweck managed to find solvable groups for which this cohomology group is non-trivial, it indicates that $G$ and $H$ are still very related to each other.
@@
The motivation to consider $FG$ lies in the fact that it is a distinguished object in $\operatorname{Rep}_{F}(G)$. Firstly, being the regular module, it is directly related to the set of irreducible $F$-representations of $G$. A second aspect is more categorical. Namely, $\operatorname{Rep}_{F}(G)$ viewed as a symmetric tensor [^1] With tensor category we mean an abelian, $F$-linear, semisimple, monoidal and rigid category whose monoidal unit is simple. category is known (by [DM, Del]) to determine completely the underlying group $G$. The latter data is equivalent to determining $FG$ as a Hopf algebra. However, solely viewed as a tensor category, there can be non-isomorphic finite groups $G$ and $H$ such that $\operatorname{Rep}_{F}(G)\simeq\operatorname{Rep}_{F}(H)$ as tensor categories, giving rise to the concept of isocategorical groups [EG, Dav, Gal]. If $F$ is algebraically closed, then $G$ and $H$ are isocategorical if and only if $FG$ and $FH$ are up to some Drinfeld twist isomorphic as Hopf algebras [EG, Proposition 3.1]. This highlights again the distinguished role of the group algebra. In our setting, one follows the natural desire to retrieve information on $G$ from much less (and more realistic to be determined) starting structure.
@@
Interestingly, if $R$ is the ring of integers of a number field, then [^2] The technical condition on $R$ allowing this conclusion, is that $R$ is a $G$-adapted ring. This means that $R$ is an integral domain of characteristic 0 such that none of the prime divisors of $|G|$ are invertible in $R$. the $R$-algebra RGRG contains the same information as its unit group ${\mathcal{U}}(RG)$. In other words, if $RG\cong RH$ as $R$-algebras, then ${\mathcal{U}}(RG)\cong{\mathcal{U}}(RH)$ as groups. Therefore, for several decades now, there has been considerable interest in understanding how $G$ is embedded in the finitely presented group ${\mathcal{U}}(RG)$. One proposed line of research is to construct a subgroup $N$ such that (1) the construction of $N$ does not depend on knowing the basis $G$ of $RG$ (i.e. it is generic in $G$), (2) it is of finite index in ${\mathcal{U}}(RG)$ and ideally (3) it is normal and torsion-free. This allows to retrieve invariants from the finite group ${\mathcal{U}}(RG)/N$. Property (2) is related with the problem of generic constructions of invertible elements [JJS].
@@
Now recall that each simple factor $\operatorname{M}_{n_{i}}(D_{i})$ in the Wedderburn–Artin decomposition (1.1) corresponds to a primitive central idempotent $e_{i}$ of $FG$. It is a general fact that ${\mathcal{U}}(RG)$ has finite index in $\prod_{i=1}^{\ell}{\mathcal{U}}(RGe_{i})$. As such, both property (2) and the normality in (3) are directly related with the description of normal subgroups of the groups ${\mathcal{U}}(RGe_{i})$. In turn, this heavily depends on the isomorphism type of $FGe_{i}=\operatorname{M}_{n_{i}}(D_{i})$. More precisely, the presence of the following type of simple algebras often breaks the standard methods used (see [BJJKT, Section 6.1] for more details):
@@
The name was coined in [MR3195747] and the reason that in practice they are exceptional (i.e. require ‘other methods’) is different. The exceptional matrix algebras are exactly those $\operatorname{M}_{n}(D)$ for which the $S$-rank of $\operatorname{SL}_{n}(D)$ is $1$, where $S$ is the set of Archimedean places of $\mathcal{Z}(D)$ (see [BJJKT, Remark 6.7] for details and eq.˜2.1 for the definition of $\operatorname{SL}_{n}$ over non-commutative rings). Consequently, they are exactly those matrix algebras having a maximal order [^3] It is well-known that a maximal order in $\operatorname{M}_{n}(D)$ is of the form $\operatorname{M}_{n}(\mathcal{O})$ with $\mathcal{O}$ a maximal order in $D$. Furthermore, the unit groups of two orders are commensurable. Thus this property does not depend on the chosen order. $\operatorname{M}_{n}(\mathcal{O})$ such that $\operatorname{SL}_{n}(\mathcal{O})$ does not satisfy the subgroup congruence problem. Moreover, there exist non-central normal subgroups which are not of finite index. The non-exceptional division algebras $D$ are exactly those having an order $\mathcal{O}$ such that $\operatorname{SL}_{1}(\mathcal{O})$ is finite. The others are problematic because $\operatorname{SL}_{1}(D)$ has no unipotent elements, which are an ingredient for most generic constructions of units in ${\mathcal{U}}(RG)$. Additionally, the subgroup congruence problem for $\operatorname{SL}_{1}(D)$ is sadly still open. Therefore, for the lines of research described earlier, irreducible representations in such simple algebras require a separate study.
@@
    Virtual cohomological dimension: for any discrete group $\Gamma$, this (denoted $\operatorname{vcd}(\Gamma)$) is the cohomological dimension of a torsion-free finite index subgroup of $\Gamma$ (see Section˜4.1 for a precise formulation).
@@
So with respect to the general theory described earlier, they represent the ‘most degenerate groups’. As a second application of understanding the properties above, we obtain several characterisations of when $FG$ has $\operatorname{(M_{exc})}$. Some of them extend the work in [FreebyFree] and others are completely new. In the sequel to this paper, we will give a precise classification of the finite groups $G$ and number fields $F$ such that $FG$ has $\operatorname{(M_{exc})}$, yielding a new case of Kleinert’s virtual structure problem [KleinertSurvey]. Moreover, we will solve the rational isomorphism problem for these groups, i.e. if ${\mathbb{Q}}G\cong{\mathbb{Q}}H$ and ${\mathbb{Q}}G$ has $\operatorname{(M_{exc})}$, then $G\cong H$.
@@ 1.1. Topological and homological characterisation of exceptional components and applications
The fact that ${\mathbb{Q}}G$ having $\operatorname{(M_{exc})}$ implies that ${\mathcal{U}}({\mathbb{Z}}G)$ is good was already obtained in [CdRZ, Theorem 1.5].
@@
In [FreebyFree] a characterisation in terms of virtual cohomological dimension and the Kleinian property was obtained of the finite groups $G$ such that ${\mathbb{Q}}G$ has no exceptional division components and the non-division components are all of the form $\operatorname{M}_{2}({\mathbb{Q}}(\sqrt{-d}))$ for some $d\in{\mathbb{N}}$. More precisely, they show that ${\mathbb{Q}}G$ has such components if and only if $\operatorname{vcd}(\operatorname{SL}_{1}({\mathbb{Z}}Ge))\leq 2$ for all $e\in\operatorname{PCI}({\mathbb{Q}}G)$. In turn the latter is equivalent to $\operatorname{SL}_{1}({\mathbb{Z}}Ge)$ having $\operatorname{Di}_{n}$ with $n\mid 2$ for all $e\in\operatorname{PCI}({\mathbb{Q}}G)$. Interestingly, including all exceptional matrix algebras adds some intricate properties on the Fitting subgroup.
In light of the aforementioned result from [FreebyFree] (and its predecessors) and ˜A, it is natural to formulate the following general problem.
@@ 1.2. Virtually free quotients and congruence kernels for higher modular groups
The first main result of Section˜5 concerns largeness for principal congruence subgroups of $\operatorname{SL}_{2}(\mathcal{O})$ with $\mathcal{O}$ an order in $D$ such that $\operatorname{M}_{2}(D)$ is an exceptional matrix algebra. A group is called large if it has a quotient which is virtually a non-abelian free group. Such groups are also referred to as having property $\operatorname{vFQ}$, see [Lubot96]. However, we will avoid this notation to avoid confusion with another property, which we will name property vQL. A group has property $\operatorname{vQL}$ if it has a finite index subgroup which maps onto a non-abelian free group, i.e. it virtually has a non-abelian free quotient. If a group is large, then it also has $\operatorname{vQL}$, since the inverse image of the free group provided by largeness is a finite index subgroup which has a free quotient. One reason for the terminology “large” is the fact that $\Gamma$ being large implies that $\Gamma$ is SQ-universal, meaning that every countable group can be embedded in a quotient of $\Gamma$.
In [Lubot96, Theorem 3.7] it was proven that if $\Gamma_{d}$ is a torsion-free finite index subgroup of $\operatorname{SL}_{2}(\mathcal{I}_{d})$, then $\Gamma_{d}$ is large. Moreover, if the prime $p$ does not split in $\mathcal{I}_{d}$, then the kernel of $\operatorname{SL}_{2}(\mathcal{I}_{d})\rightarrow\operatorname{SL}_{2}(\mathcal{I}_{d}/p\mathcal{I}_{d})$ maps onto a free group of rank $p^{3}+p-1$. Note that the property $\operatorname{vQL}$ is stable under commensurability. Hence [Lubot96, Theorem 3.7] gave a constructive version of Grunewald–Schwermer’s theorem [GS], stating that $\operatorname{SL}_{2}(\mathcal{I}_{d})$ has $\operatorname{vQL}$. We obtain the analogous statement for $\operatorname{SL}_{2}(\mathcal{O})$ with $\mathcal{O}$ an order in $\left(\frac{u,v}{{\mathbb{Q}}}\right)$, where $u$ and $v$ are strictly negative integers.
@@
The proof of ˜B is geometric and uses tools from [Lubot96] and [Kiefer]. More precisely, in Section˜5.2 we relate the principal congruence subgroups of $\operatorname{SL}_{2}(\mathcal{L}_{u,v})$ to congruence subgroups of certain groups of isometries of the hyperbolic $5$-space, denoted $\operatorname{SL}_{+}\left(\Gamma_{4}^{u,v}({\mathbb{Z}})\right)$. To do so, in Section˜5.1 we recall how studying $2\times 2$ matrix groups over orders in quaternion algebras all comes down to studying discrete subgroups $\operatorname{SL}_{+}\left(\Gamma_{4}({\mathbb{Z}})\right)$ of Vahlen groups, which are groups of $2\times 2$ matrices with entries in a Clifford algebra. These groups were described in [Vahlen], where Vahlen generalised Möbius transformations to higher-dimensional hyperbolic spaces. Discrete subgroups of the latter provide a natural generalisation of the modular group to higher dimensions. These generalised groups are denoted by $\operatorname{SL}_{+}(\Gamma_{n}({\mathbb{Z}}))$ and referred to as higher modular groups. ˜B will be a consequence of the following statement on higher modular groups. In fact, the following was our motivation to consider largeness of congruence subgroups of $\operatorname{SL}_{2}(\mathcal{L}_{u,v})$.
@@
Being large implies that the abelianisation is infinite. In particular, if a group $\Gamma$ has $\operatorname{vQL}$, then $\Gamma$ does not have property (FAb), i.e. not every finite index subgroup has finite abelianisation. This observation, along with Theorem˜5.1 applied to $\left(\frac{-2,-5}{{\mathbb{Q}}}\right)$, can be used to remove the condition $5\nmid|G|$ in [BJJKT2, Theorem 6.10].
@@
In Section˜5.3 we use ˜D (resp. ˜B) to show that the congruence kernel of $\operatorname{SL}_{+}\left(\Gamma_{4}^{u,v}({\mathbb{Z}})\right)$ (resp. $\operatorname{SL}_{1}(\mathcal{O})$) contains as a closed subgroup a copy of the free group of countable rank, see Theorem˜5.7 and Corollary˜5.8. In case of $\operatorname{SL}_{2}(\mathcal{I}_{d})$ this was obtained in [Lubot82, Theorem B] and our proof uses the methods in loc. cit.

### v2

@@ 1. Introduction
This problem is highly dependent on the choice of the ground field $F$. For example if $F={\mathbb{C}}$, then $FG\cong FH$ for $G$ and $H$ abelian exactly when $|G|=|H|$. On the other hand, if $F={\mathbb{Q}}$ then the group algebra determines all the structure constants of an abelian group by a theorem of Perlis–Walker [70]. One reason that $\mathbb{Q}$} detects more is the theorem of Wedderburn–Artin, which says that there is an isomorphism of $F$-algebras
@@
where $D_{i}$ are finite-dimensional division $F$-algebras. In particular, the Brauer group ${\rm Br}(F)$, i.e. the division algebras over $F$, has an influence. One way to view this is, the ‘smaller’ $F$, the ‘richer’ its Brauer group. Nevertheless, Dade [16] constructed non-isomorphic groups $G$ and $H$ with isomorphic group algebras over any field of characteristic 0, or in other words such that ${\mathbb{Q}}G\cong{\mathbb{Q}}H$. For a long time, it was believed that over the integers $\mathbb{Z}$} (equivalently over all group rings $RG$), a group $G$ is uniquely determined by its integral group ring ${\mathbb{Z}}G$. However, Hertweck constructed a counterexample to this [37]. Nevertheless, if ${\mathbb{Z}}G\cong{\mathbb{Z}}H$, then $G$ and $H$ are cocycle twists of each under a Cech-type non-abelian cohomology theory constructed by Roggenkamp–Kimmerle [53]. Thus, although Hertweck managed to find solvable groups for which this cohomology group is non-trivial, it indicates that $G$ and $H$ are still very related to each other.
@@
The motivation to consider $FG$ lies in the fact that it is a distinguished object in $\operatorname{Rep}_{F}(G)$. Firstly, being the regular module, it is directly related to the set of irreducible $F$-representations of $G$. A second aspect is more categorical. Namely, $\operatorname{Rep}_{F}(G)$ viewed as a symmetric tensor [^1] With tensor category we mean an abelian, $F$-linear, semisimple, monoidal and rigid category whose monoidal unit is simple. category is known (by [20, 19]) to determine completely the underlying group $G$. The latter data is equivalent to determining $FG$ as a Hopf algebra. However, solely viewed as a tensor category, there can be non-isomorphic finite groups $G$ and $H$ such that $\operatorname{Rep}_{F}(G)\simeq\operatorname{Rep}_{F}(H)$ as tensor categories, giving rise to the concept of isocategorical groups [26, 17, 28]. If $F$ is algebraically closed, then $G$ and $H$ are isocategorical if and only if $FG$ and $FH$ are up to some Drinfeld twist isomorphic as Hopf algebras [26, Proposition 3.1]. This highlights again the distinguished role of the group algebra. In our setting, one follows the natural desire to retrieve information on $G$ from much less (and more realistic to be determined) starting structure.
@@
Interestingly, if $R$ is the ring of integers of a number field, then [^2] The technical condition on $R$ allowing this conclusion, is that $R$ is a $G$-adapted ring. This means that $R$ is an integral domain of characteristic 0 such that none of the prime divisors of $|G|$ are invertible in $R$. the $R$-algebra RGRG contains the same information as its unit group ${\mathcal{U}}(RG)$. In other words, if $RG\cong RH$ as $R$-algebras, then ${\mathcal{U}}(RG)\cong{\mathcal{U}}(RH)$ as groups. Therefore, for several decades now, there has been considerable interest in understanding how $G$ is embedded in the finitely presented group ${\mathcal{U}}(RG)$. One proposed line of research is to construct a subgroup $N$ such that (1) the construction of $N$ does not depend on knowing the basis $G$ of $RG$ (i.e. it is generic in $G$), (2) it is of finite index in ${\mathcal{U}}(RG)$ and ideally (3) it is normal and torsion-free. This allows to retrieve invariants from the finite group ${\mathcal{U}}(RG)/N$. Property (2) is related with the problem of generic constructions of invertible elements [43].
@@
Now recall that each simple factor $\operatorname{M}_{n_{i}}(D_{i})$ in the Wedderburn–Artin decomposition (1.1) corresponds to a primitive central idempotent $e_{i}$ of $FG$; we denote the set of all such elements by $\operatorname{PCI}(FG)$. It is a general fact that ${\mathcal{U}}(RG)$ has finite index in $\prod_{i=1}^{\ell}{\mathcal{U}}(RGe_{i})$. As such, both property (2) and the normality in (3) are directly related with the description of normal subgroups of the groups ${\mathcal{U}}(RGe_{i})$. In turn, this heavily depends on the isomorphism type of $FGe_{i}=\operatorname{M}_{n_{i}}(D_{i})$. More precisely, the presence of the following type of simple algebras often breaks the standard methods used (see [4, Section 6.1] for more details):
@@
The name was coined in [49] and the reason that in practice they are exceptional (i.e. require ‘other methods’) is different. The exceptional matrix algebras are exactly those $\operatorname{M}_{n}(D)$ for which the $S$-rank of $\operatorname{SL}_{n}(D)$ is $1$, where $S$ is the set of Archimedean places of $\mathcal{Z}(D)$ (see [4, Remark 6.7] for details and eq.˜2.1 for the definition of $\operatorname{SL}_{n}$ over non-commutative rings). Consequently, they are exactly those matrix algebras having a maximal order [^3] It is well-known that a maximal order in $\operatorname{M}_{n}(D)$ is of the form $\operatorname{M}_{n}(\mathcal{O})$ with $\mathcal{O}$ a maximal order in $D$. Furthermore, the unit groups of two orders are commensurable. Thus this property does not depend on the chosen order. $\operatorname{M}_{n}(\mathcal{O})$ such that $\operatorname{SL}_{n}(\mathcal{O})$ does not satisfy the subgroup congruence problem. Moreover, there exist non-central normal subgroups which are not of finite index. The non-exceptional division algebras $D$ are exactly those having an order $\mathcal{O}$ such that $\operatorname{SL}_{1}(\mathcal{O})$ is finite. The others are problematic because $\operatorname{SL}_{1}(D)$ has no unipotent elements, which are an ingredient for most generic constructions of units in ${\mathcal{U}}(RG)$. Additionally, the subgroup congruence problem for $\operatorname{SL}_{1}(D)$ is sadly still open. Therefore, for the lines of research described earlier, irreducible representations in such simple algebras require a separate study.
@@
    Virtual cohomological dimension: for any discrete group $\Gamma$, this is the cohomological dimension of a torsion-free finite index subgroup of $\Gamma$, and we denote it by $\operatorname{vcd}(\Gamma)$ (see Section˜4.1 for a precise formulation).
@@
So with respect to the general theory described earlier, they represent the ‘most degenerate groups’. As a second application of understanding the properties above, we obtain several characterisations of when $FG$ has $\operatorname{(M_{exc})}$. Some of them extend the work in [50] and others are completely new. In the sequel to this paper, we will give a precise classification of the finite groups $G$ and number fields $F$ such that $FG$ has $\operatorname{(M_{exc})}$, yielding a new case of Kleinert’s virtual structure problem [56]. Moreover, we will solve the rational isomorphism problem for these groups, i.e. if ${\mathbb{Q}}G\cong{\mathbb{Q}}H$ and ${\mathbb{Q}}G$ has $\operatorname{(M_{exc})}$, then $G\cong H$.
@@ 1.1. Topological and homological characterisation of exceptional components and applications
The fact that ${\mathbb{Q}}G$ having $\operatorname{(M_{exc})}$ implies that ${\mathcal{U}}({\mathbb{Z}}G)$ is good was already obtained in [13, Theorem 1.5].
@@
In [50] a characterisation in terms of virtual cohomological dimension and the Kleinian property was obtained of the finite groups $G$ such that ${\mathbb{Q}}G$ has no exceptional division components and the non-division components are all of the form $\operatorname{M}_{2}({\mathbb{Q}}(\sqrt{-d}))$ for some $d\in{\mathbb{N}}$. More precisely, they show that ${\mathbb{Q}}G$ has such components if and only if $\operatorname{vcd}(\operatorname{SL}_{1}({\mathbb{Z}}Ge))\leq 2$ for all $e\in\operatorname{PCI}({\mathbb{Q}}G)$. In turn the latter is equivalent to $\operatorname{SL}_{1}({\mathbb{Z}}Ge)$ having $\operatorname{Di}_{n}$ with $n\mid 2$ for all $e\in\operatorname{PCI}({\mathbb{Q}}G)$. Interestingly, including all exceptional matrix algebras adds some intricate properties on the Fitting subgroup.
In light of the aforementioned result from [50] (and its predecessors) and ˜A, it is natural to formulate the following general problem.
@@ 1.2. Virtually free quotients and congruence kernels for higher modular groups
The first main result of Section˜5 concerns largeness for principal congruence subgroups of $\operatorname{SL}_{2}(\mathcal{O})$ with $\mathcal{O}$ an order in $D$ such that $\operatorname{M}_{2}(D)$ is an exceptional matrix algebra. A group is called large if it has a quotient which is virtually a non-abelian free group. Such groups are also referred to as having property $\operatorname{vFQ}$, see [59]. However, we will avoid this notation to avoid confusion with another property, which we will name property vQL. A group has property $\operatorname{vQL}$ if it has a finite index subgroup which maps onto a non-abelian free group, i.e. it virtually has a non-abelian free quotient. If a group is large, then it also has $\operatorname{vQL}$, since the inverse image of the free group provided by largeness is a finite index subgroup which has a free quotient. One reason for the terminology “large” is the fact that $\Gamma$ being large implies that $\Gamma$ is SQ-universal, meaning that every countable group can be embedded in a quotient of $\Gamma$.
In [59, Theorem 3.7] it was proven that if $\Gamma_{d}$ is a torsion-free finite index subgroup of $\operatorname{SL}_{2}(\mathcal{I}_{d})$, then $\Gamma_{d}$ is large. Moreover, if the prime $p$ does not split in $\mathcal{I}_{d}$, then the kernel of $\operatorname{SL}_{2}(\mathcal{I}_{d})\rightarrow\operatorname{SL}_{2}(\mathcal{I}_{d}/p\mathcal{I}_{d})$ maps onto a free group of rank $p^{3}+p-1$. Note that the property $\operatorname{vQL}$ is stable under commensurability. Hence [59, Theorem 3.7] gave a constructive version of Grunewald–Schwermer’s theorem [32], stating that $\operatorname{SL}_{2}(\mathcal{I}_{d})$ has $\operatorname{vQL}$. We obtain the analogous statement for $\operatorname{SL}_{2}(\mathcal{O})$ with $\mathcal{O}$ an order in $\left(\frac{u,v}{{\mathbb{Q}}}\right)$, where $u$ and $v$ are strictly negative integers.
@@
The proof of ˜B is geometric and uses tools from [59] and [52]. More precisely, in Section˜5.2 we relate the principal congruence subgroups of $\operatorname{SL}_{2}(\mathcal{L}_{u,v})$ to congruence subgroups of certain groups of isometries of the hyperbolic $5$-space, denoted $\operatorname{SL}_{+}\left(\Gamma_{4}^{u,v}({\mathbb{Z}})\right)$. To do so, in Section˜5.1 we recall how studying $2\times 2$ matrix groups over orders in quaternion algebras all comes down to studying discrete subgroups $\operatorname{SL}_{+}\left(\Gamma_{4}({\mathbb{Z}})\right)$ of Vahlen groups, which are groups of $2\times 2$ matrices with entries in a Clifford algebra. These groups were described in [84], where Vahlen generalised Möbius transformations to higher-dimensional hyperbolic spaces. Discrete subgroups of the latter provide a natural generalisation of the modular group to higher dimensions. These generalised groups are denoted by $\operatorname{SL}_{+}(\Gamma_{n}({\mathbb{Z}}))$ and referred to as higher modular groups. ˜B will be a consequence of the following statement on higher modular groups. In fact, the following was our motivation to consider largeness of congruence subgroups of $\operatorname{SL}_{2}(\mathcal{L}_{u,v})$.
@@
Being large implies that the abelianisation is infinite. In particular, if a group $\Gamma$ has $\operatorname{vQL}$, then $\Gamma$ does not have property (FAb), i.e. not every finite index subgroup has finite abelianisation. This observation, along with Theorem˜5.1 applied to $\left(\frac{-2,-5}{{\mathbb{Q}}}\right)$, can be used to remove the condition $5\nmid|G|$ in [3, Theorem 6.10].
@@
In Section˜5.3 we use ˜D (resp. ˜B) to show that the congruence kernel of $\operatorname{SL}_{+}\left(\Gamma_{4}^{u,v}({\mathbb{Z}})\right)$ (resp. $\operatorname{SL}_{1}(\mathcal{O})$) contains as a closed subgroup a copy of the free group of countable rank, see Theorem˜5.7 and Corollary˜5.8. In case of $\operatorname{SL}_{2}(\mathcal{I}_{d})$ this was obtained in [58, Theorem B] and our proof uses the methods in loc. cit.